OpenStudy (anonymous):

solve the system using substation 3a-b=11 2a+3b= 0

OpenStudy (anonymous):

Manipulate the first algebraic equation to get b on one side of the equation. You should get an equation that looks like 3a-11=b. Next thing you do is then plug this equation into the second equation and you get 2a+3(3a-11) = 0. Distribute this out to get 2a+9a-33=0. Solve for a. If you do it correctly you should get a = 3.

OpenStudy (anonymous):

what about solving for b?

OpenStudy (anonymous):

Once you get a which is a =3 you can then 3 as a into either one of the equations. You should get b = -2 when you do this. You can plug (3) as a into both equations and you should get b = -2 for both of the equations.

OpenStudy (anonymous):

oh! okay.

OpenStudy (anonymous):

So for example 3a - 11 = b. 3(3) - 11 = b. 9 -11 = b. -2 = b. Other equation 2(3) + 3b = 0. 6 + 3b = 0 3b = -6 b = -2